Check whether the set of vectors {1 + π‘₯, π‘₯ + π‘₯2


, 1 + π‘₯3


} is a linearly independent set of


vectors in P3


, the vector space of polynomials of degree ≀ 3

1
Expert's answer
2022-04-11T07:36:57-0400

Suppose that there are complex numbers "\alpha_1,\alpha_2,\alpha_3" that are not all equal to zero and satisfy the equality: "\alpha_1(1+x)+\alpha_2(1+x^2)+\alpha_3(1+x^3)=0" for all "x". We came to contradiction because a cubic equation may have only finite number of roots. Another way to come to a contradiction is to substitute values "x=0,x=2,x=3." We receive equations "\alpha_1+\alpha_2+\alpha_3=0", "3\alpha_1+5\alpha_2+9\alpha_3=0", "4\alpha_1+10\alpha_2+28\alpha_3=0". Since "\left|\begin{array}{lll}1&1&1\\3&5&9\\4&10&28\end{array}\right|=28\neq0", the latter system of equations has the unique solution "\alpha_1=\alpha_2=\alpha_3=0". Thus, due to the deduced contradiction, vectors "1+x,1+x^2,1+x^3" are linearly independent.